# Style-NeRF2NeRF: 3D Style Transfer from Style-Aligned Multi-View Images

### v2

@@ Abstract
Keywords Neural Radiance Fields  $\cdot$ Style Transfer  $\cdot$ Sliced Wasserstein
1 Introduction
Thanks to recent advancements in 3D reconstruction techniques such as Neural Radiance Fields (NeRF) Mildenhall et al. (2020), it is nowadays possible for creators to develop a 3D asset or a scene from captured real-world data without intensive labor. While such 3D reconstruction methods work well, editing an entire 3D scene to match a desired style or concept is not straightforward.
@@ 1 Introduction
At the advent of implicit 3D representation techniques such as NeRF, style editing methods for 3D are also emerging Nguyen-Phuoc et al. (2022); Wang et al. (2023); Liu et al. (2023); Kamata et al. (2023); Haque et al. (2023); Dong and Wang (2024) to enhance creators’ content development process. Following the recent development of 2D image generation models, prominent works such as Instruct-NeRF2NeRF Haque et al. (2023); Vachha and Haque (2024) and ViCA-NeRF Dong and Wang (2024) proposed to leverage the knowledge of large-scale pre-trained text-to-image (T2I) models to supervise the 3D NeRF editing process.
These methods employ a custom pipeline based on an instruction-based T2I model ”Instruct-Pix2Pix” Brooks et al. (2023) to stylize a 3D scene with text instructions. While Instruct-NeRF2NeRF is proven to work well for editing 3D scenes including large-scale 360 environments, their method involves an iterative process of editing and replacing the training data during NeRF optimization, occasionally resulting in unpredictable results. As editing by Instruct-Pix2Pix runs in tandem with NeRF training, we found adjusting or testing editing styles beforehand difficult.
To overcome this problem, we propose an artistic style-transfer method that trains a source 3D NeRF scene on stylized images prepared in advance by a text-guided style-aligned diffusion model. Training is guided by Sliced Wasserstein Distance (SWD) loss Heitz et al. (2021); Li et al. (2022) to effectively perform 3D style transfer with NeRF. A summary of our contributions is as the follows:
@@
2 Related Work
2.1 Implicit 3D Representation
@@
NeRF, introduced by the seminal paper Mildenhall et al. (2020) became one of the most popular implicit 3D representation techniques due to several benefits. NeRF can render photo-realistic novel views with arbitrary resolution due to its continuous representation with a compact model compared to explicit representations such as polygon mesh or voxels. In our research, we use the "nerfacto" model implemented by Nerfstudio Tancik et al. (2023), which is a combination of modular features from multiple papers Wang et al. (2021); Barron et al. (2022); Müller et al. (2022); Martin-Brualla et al. (2021); Verbin et al. (2022) , designed to achieve a balance between speed and quality.
2.2 Style Transfer
2.2.1 2D Style Transfer.
Style transfer is a technique for blending images, a source image and a style image, to create another image that retains the first’s content but exhibits the second’s style. Since the introduction of the foundational style transfer algorithm proposed by Gatys et al. (2015), many follow-up works for 2D style transfer have been explored for further improvements such as faster optimization Johnson et al. (2016); Huang and Belongie (2017), zero-shot style-transfer Li et al. (2017), and photo-realism Luan et al. (2017).
2.2.2 3D Style Transfer.
Several recent 3D style transfer works have applied style transfer techniques using deep feature statistics Huang and Belongie (2017) to NeRF Liu et al. (2023); Wang et al. (2023); Zhang et al. (2022). Although these methods require a reference style image, text-based 3D editing techniques have also been proposed leveraging foundational 2D Text-to-Image (T2I) generative models. While Instruct 3D-to-3D Kamata et al. (2023) proposed using Score Distillation Sampling (SDS) loss Poole et al. (2022) for text guided NeRF stylization, Instruct-NeRF2NeRF Haque et al. (2023) and ViCA-NeRF Dong and Wang (2024) perform NeRF editing by optimizing the underlying scene with a process referred to as Iterative Dataset Update (Iterative DU), which gradually replaces the input images with edited images from InstructPix2Pix Haque et al. (2023), an image-conditioned instruction-based diffusion model, followed by an update of NeRF. Inspired by these methods, we also develop a 3D style transfer method for NeRF, supervised by images created by a diffusion pipeline but without Iterative DU.
2.3 Diffusion Models
@@
Diffusion models Sohl-Dickstein et al. (2015); Song et al. (2020a); Dhariwal and Nichol (2021) are generative models that have gained significant attention for their ability to generate high-quality, diverse images. Inspired by classical non-equilibrium thermodynamics, they are trained to generate an image by reversing the diffusion process, progressively denoising noisy images towards meaningful ones. Diffusion models are commonly trained with classifier-free guidance Ho and Salimans (2022) to enable image generation conditioned on an input text.
2.3.1 Controlled Generations with Diffusion Models.
@@
Leveraging the success of T2I diffusion models, recent research has expanded their application to controlled image generation and editing, notably in image-to-image (I2I) tasks Meng et al. (2021); Parmar et al. (2023); Kawar et al. (2023); Tumanyan et al. (2023); Mokady et al. (2023); Hertz et al. (2023a, 2022); Brooks et al. (2023). For example, SDEdit Meng et al. (2021) achieves this by first adding noise to a source image and then guiding the diffusion process toward an output based on a given prompt. ControlNet Zhang et al. (2023) was proposed as an add-on architecture for training T2I diffusion models with extra conditioning inputs such as depth, pose, edge maps, and more. Several recent techniques Hertz et al. (2023b); Sohn et al. (2024); Cheng et al. (2023) focus on generating style-aligned images. In our work, we use a depth-conditioned I2I pipeline with an attention-sharing mechanism similar to "StyleAligned" Hertz et al. (2023b) to create a set of multi-view images sharing a consistent style.
@@
Figure 1: Overall Pipeline: Our method consists of distinct procedures. We first prepare a NeRF model of the source view images. Given the depth maps of the corresponding views (by either estimation or rendering by NeRF), we generate stylized multi-view images using a style-aligned diffusion model. Lastly, we fine-tune the source NeRF on the stylized images using the SWD loss.
3 Method
3.1 Preliminaries
3.1.1 Neural Radiance Fields.
@@ 3.1 Preliminaries
NeRF Mildenhall et al. (2020) models a volumetric 3D scene as a continuous function by mapping a 3D coordinate $\mathbf{x}=(x,y,z)$ and a 2D viewing direction $\mathbf{d}=(\theta,\phi)$ to a color (RGB) $\mathbf{c}$ and a density ($\sigma$). This function $F_{\theta}:(\mathbf{x},\mathbf{d})\mapsto(\mathbf{c},\mathbf{\sigma})$ is often parameterized by a neural network combined with voxel grid structures or other encoding techniques Fridovich-Keil et al. (2022); Müller et al. (2022); Sun et al. (2022a, b) to accelerate performance. Given a NeRF model trained on a set of 2D images taken from various viewpoints of a target scene, the accumulated color $C(\mathbf{r})$ along an arbitrary camera ray $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ is calculated with the quadrature rule by volume rendering Max (1995):
$C(\mathbf{r})=\sum_{k=1}^{K}T_{k}(1-\exp{-\sigma_{k}\delta_{k}})\mathbf{c}_{k},T_{k}=\exp{\Bigl{(}-\sum_{j=1}^{k-1}\sigma_{j}\delta_{j}\Bigr{)}}$ (1)
@@
3.1.2 Conditional Diffusion Models.
Recent T2I diffusion models Rombach et al. (2022); Podell et al. (2023) are built with a U-net architecture Ronneberger et al. (2015) integrated with convolutional layers and attention blocks Vaswani et al. (2017). Within the model, attention blocks play a crucial role in correlating text with relevant parts of the deep features during image generation. Our work uses an open-source latent diffusion model Podell et al. (2023), which includes a CLIP text encoder Radford et al. (2021) for text embedding. The cross-attention between contextual text embedding and the deep features of the denoising network is calculated as follows:
$\texttt{Attn}(Q,K,V)=\texttt{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}V\right)$ (2)
@@
$\texttt{MultiHeadAttn}(Q,K,V)=\underset{j\in N_{h}}{\texttt{concat}}\Bigl{[}\texttt{Attn}(Q^{(j)},K^{(j)},V^{(j)})\Bigr{]}$ (3)
3.2 Style-NeRF2NeRF
@@
3.2.1 Style-Aligned Image-to-Image Generation.
Given a set of source view images $\{I^{(i)}\}(i=1,\ldots,N)$, our first goal is to generate a corresponding set of stylized view images $I^{(i)}_{c}=U_{\theta}(I^{(i)},c)$ under a text condition $c$ with as much perceptual view consistencies among images where $U_{\theta}$ consists of a sampling process such as DDIM Song et al. (2020b).
Although T2I diffusion models can generate rich images with arbitrary text prompts, merely sharing the same prompt across different source views is insufficient to generate stylized images with a perceptually consistent style. To alleviate this problem, we apply a fully-shared-attention variant of a style-aligned image generation method proposed by Hertz et al. (2023b). Let $Q_{i},K_{i},V_{i}$ be the queries, keys, and values from a deep feature $\phi_{i}$ for view $I^{(i)}$, then we generate $n$ stylized views simultaneously using the following fully-shared-attention:
$\texttt{Attn}(Q_{i},K_{1\ldots n},V_{1\ldots n})\$ (4)
$K_{1\ldots n}=[K_{1},K_{2},\ldots K_{n}]^{T},V_{1\ldots n}=[V_{1},V_{2},\ldots V_{n}]^{T}$ (5)
@@
3.2.2 Conditioning on Source Views.
To further strengthen perceptual consistencies across multi-view frames, we attach a depth-conditioned ControlNet Zhang et al. (2023) and optionally enable SDEdit Meng et al. (2021) for conditioning on the source view. As for the depth inputs, we may either render the corresponding depth maps from the source NeRF or use an off-the-shelf depth estimator model such as MiDaS Ranftl et al. (2020).
@@
3.2.3 NeRF Fine-Tuning.
@@
Although the stylized multi-view images are a good starting point for fine-tuning the source NeRF, we found that using a common RGB pixel loss is prone to over-fitting due to ambiguities in 3D geometry and color. Therefore, an alternative loss function that reflects the perceptual similarity is preferred for guiding the 3D style-transfer process. To meet our requirement, we employ the Sliced Wasserstein Distance loss (SWD loss) Heitz et al. (2021).
3.3 Sliced Wasserstein Distance Loss.
Feature statistics of pre-trained Convolutional Neural Networks (CNNs) such as VGG-19 Simonyan and Zisserman (2014) are known to be useful for representing a style of an image Gatys et al. (2015); Johnson et al. (2016); Huang and Belongie (2017); Li et al. (2017); Luan et al. (2017). In our study we employ the SWD loss originally proposed for texture synthesis Heitz et al. (2021) as the loss term to guide the style-transfer process for NeRF.
@@
$\mathcal{L}_{style}=\sum_{l=1}^{L}\mathcal{L}_{SWD}(p^{l},\hat{p}^{l})$ (7)
@@
$\mathcal{L}_{SWD}=\sum_{l=1}^{L}\mathbb{E}_{V}[\mathcal{L}_{SW1D}(p^{l}_{V},\hat{p}^{l}_{V})]$ (8)
$\mathcal{L}_{SW1D}(p^{l}_{V},\hat{p}^{l}_{V})=\frac{1}{|p^{l}_{V}|}\lVert\texttt{sort}(p^{l}_{V})-\texttt{sort}(\hat{p}^{l}_{V})\rVert^{2}$ (9)
Expectation over random projections $V$ provides a good approximation in practice and an optimized distribution is proven to converge to the target distribution. SWD is known to capture the complete target distribution Pitie et al. (2005) as described below:
$\mathcal{L}_{SW}(I,\hat{I})=0\implies p^{l}=\hat{p}^{l},\forall l\in{1,\ldots,L}$ (10)
@@
Figure 2: Sliced Wasserstein Distance: $p$ and $\hat{p}$ are projected onto a random unit direction $V$ (left). The $1$-dimensional Wasserstein distance can be calculated by taking the $L^{2}$ difference between the sorted projections $p$ and $\hat{p}$ (right). Expectation over random $V$ vectors is a practical approximation of the $N$-dimensional Wasserstein distance.
3.4 Style Blending.
@@
$\mathcal{L}_{style}(I_{1},I_{2},\hat{I})=\sum_{l=1}^{L}\left(t\mathcal{L}_{SWD}(p_{1}^{l},\hat{p}^{l})+(1-t)\mathcal{L}_{SWD}(p_{2}^{l},\hat{p}^{l})\right)$ (11)
@@
Figure 3: Style Interpolation: An example of style blending using the Wasserstein barycenter between two different style prompts "A person like Marilyn Monroe, pop art style" and "A person like Steve Jobs".
3.5 Implementation Details.
@@
We employ Stable Diffusion XL Podell et al. (2023) as a backbone for the style-aligned image-to-image diffusion pipeline. For NeRF representation, we use the "nerfacto" model implemented in NeRFStudio Tancik et al. (2023). Due to memory constraints, we generate up to $18$ views simultaneously with a fixed seed across all $N$ source views. In our experiments, we generated the target images with $50$ denoising steps using a range of classifier-free guidance weights (mostly between 5 and 30) depending on the scene or the style text. While we use depth maps rendered by NeRF for relatively compact and forward-facing scenes, we opt for depth estimations from the MiDaS model Ranftl et al. (2020) for large-scale outdoor scenes.
@@
As the image editing is performed before NeRF training, our method allows users to test with different text prompts and parameters (e.g., text guidance scale, SDEdit strength) beforehand. Additionally, our straight-forward NeRF training without Iterative DU Haque et al. (2023) or score distillation sampling (SDS) loss Poole et al. (2022) allows the training process to run with less GPU memory as editing by a diffusion model is not necessary during NeRF updates. We also verify the importance of style-alignment in the ablation study. Please refer to the supplementary material for more implementation details. The overall pipeline of our method is shown in Figure 1.
@@
Figure 4: Effect of Style-Alignment: An example of source view conversion applied to "Bear" scene using a text prompt "A bear, impressionism painting style" with and without shared-attention mechanism within the diffusion pipeline. We find that a fully-shared-attention variant of the style-aligned diffusion model Hertz et al. (2023b) greatly improves perceptual consistencies among generated views.
4 Results

### v3

@@ Abstract.
Neural Radiance Fields, Neural Rendering, Style Transfer, Diffusion Model
ccs: Computing methodologies Non-photorealistic renderingccs: Computing methodologies Computer vision representations
Figure 1. Our method makes it possible to perform 3D artistic style transfer on a pre-trained NeRF scene using text descriptions.
1. Introduction
Thanks to recent advancements in 3D reconstruction techniques such as Neural Radiance Fields (NeRF) (Mildenhall et al., 2020), it is nowadays possible for creators to develop a 3D asset or a scene from captured real-world data without intensive labor. While such 3D reconstruction methods work well, editing an entire 3D scene to match a desired style or concept is not straightforward.
@@ 1. Introduction
At the advent of implicit 3D representation techniques such as NeRF, style editing methods for 3D are also emerging (Nguyen-Phuoc et al., 2022; Wang et al., 2023; Liu et al., 2023; Kamata et al., 2023; Haque et al., 2023; Dong and Wang, 2024) to enhance creators’ content development process. Following the recent development of 2D image generation models, prominent works such as Instruct-NeRF2NeRF (Haque et al., 2023; Vachha and Haque, 2024) and ViCA-NeRF (Dong and Wang, 2024) proposed to leverage the knowledge of large-scale pre-trained text-to-image (T2I) models to supervise the 3D NeRF editing process.
These methods employ a custom pipeline based on an instruction-based T2I model ”Instruct-Pix2Pix” (Brooks et al., 2023) to stylize a 3D scene with text instructions. While Instruct-NeRF2NeRF is proven to work well for editing 3D scenes including large-scale 360 environments, their method involves an iterative process of editing and replacing the training data during NeRF optimization, occasionally resulting in unpredictable results. As editing by Instruct-Pix2Pix runs in tandem with NeRF training, we found adjusting or testing editing styles beforehand difficult.
To overcome this problem, we propose an artistic style-transfer method that trains a source 3D NeRF scene on stylized images prepared in advance by a text-guided style-aligned diffusion model. Training is guided by Sliced Wasserstein Distance (SWD) loss (Heitz et al., 2021; Li et al., 2022) to effectively perform 3D style transfer with NeRF. A summary of our contributions is as the follows:
@@
2. Related Work
2.1. Implicit 3D Representation
@@
NeRF, introduced by the seminal paper (Mildenhall et al., 2020), became one of the most popular implicit 3D representation techniques due to several benefits. NeRF can render photo-realistic novel views with arbitrary resolution due to its continuous representation with a compact model compared to explicit representations such as polygon mesh or voxels. In our research, we use the ”nerfacto” model implemented by Nerfstudio (Tancik et al., 2023), which is a combination of modular features from multiple papers (Wang et al., 2021; Barron et al., 2022; Müller et al., 2022; Martin-Brualla et al., 2021; Verbin et al., 2022) , designed to achieve a balance between speed and quality.
2.2. Diffusion Models
@@
Diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2020b; Dhariwal and Nichol, 2021) are generative models that have gained significant attention for their ability to generate high-quality, diverse images. Inspired by classical non-equilibrium thermodynamics, they are trained to generate an image by reversing the diffusion process, progressively denoising noisy images towards meaningful ones. Diffusion models are commonly trained with classifier-free guidance (Ho and Salimans, 2022) to enable image generation conditioned on an input text.
2.2.1. Controlled Generations with Diffusion Models.
@@
Leveraging the success of T2I diffusion models, recent research has expanded their application to controlled image generation and editing, notably in image-to-image (I2I) tasks (Meng et al., 2021; Parmar et al., 2023; Kawar et al., 2023; Tumanyan et al., 2023; Mokady et al., 2023; Hertz et al., 2023a, 2022; Brooks et al., 2023). For example, SDEdit (Meng et al., 2021) achieves this by first adding noise to a source image and then guiding the diffusion process toward an output based on a given prompt. ControlNet (Zhang et al., 2023) was proposed as an add-on architecture for training T2I diffusion models with extra conditioning inputs such as depth, pose, edge maps, and more. Several recent techniques (Hertz et al., 2023b; Sohn et al., 2024; Cheng et al., 2023) focus on generating style-aligned images. In our work, we use a depth-conditioned I2I pipeline with an attention-sharing mechanism similar to ”StyleAligned” (Hertz et al., 2023b) to create a set of multi-view images sharing a consistent style.
2.3. Style Transfer
2.3.1. 2D Style Transfer.
Style transfer originally refers to a technique for blending images, a source image and a style image, to create another image that retains the first’s content but exhibits the second’s style. Since the introduction of the foundational style transfer algorithm proposed by (Gatys et al., 2015), many follow-up works for 2D style transfer have been explored for further improvements such as faster optimization (Johnson et al., 2016; Huang and Belongie, 2017), zero-shot style-transfer (Li et al., 2017), and photo-realism (Luan et al., 2017). Furthermore, content stylization methods using only text descriptions for style (Frenkel et al., 2024; Sohn et al., 2024; Shah et al., 2023) are showing promising results due to the recent progress in controllable diffusion models.
2.3.2. 3D Style Transfer.
Several recent 3D style transfer works have applied style transfer techniques using deep feature statistics to NeRF (Liu et al., 2023; Wang et al., 2023; Zhang et al., 2022; Chiang et al., 2022; Huang et al., 2022; Nguyen-Phuoc et al., 2022; Pang et al., 2023). In addition to such stylization methods based on a style reference, text-driven 3D editing techniques leveraging foundational 2D Text-to-Image (T2I) models are developed. While Instruct 3D-to-3D (Kamata et al., 2023) proposed using Score Distillation Sampling (SDS) loss (Poole et al., 2022) for text guided NeRF stylization, Instruct-NeRF2NeRF (Haque et al., 2023) and ViCA-NeRF (Dong and Wang, 2024) perform NeRF editing by optimizing the underlying scene with a process referred to as Iterative Dataset Update (Iterative DU), which gradually replaces the input images with edited images from InstructPix2Pix (Haque et al., 2023), an image-conditioned instruction-based diffusion model, followed by an update of NeRF. Inspired by these methods, we also develop a 3D style transfer method for NeRF, supervised by images created by a diffusion pipeline but without Iterative DU.
@@ 2.3. Style Transfer
Figure 2. Overall Pipeline. Our method consists of distinct procedures. We first prepare a NeRF model of the source view images. Given the depth maps of the corresponding views (by either estimation or rendering by NeRF), we generate stylized multi-view images using a style-aligned diffusion model. Lastly, we fine-tune the source NeRF on the stylized images using the SWD loss.
3. Method
3.1. Preliminaries
3.1.1. Neural Radiance Fields.
@@ 3.1. Preliminaries
NeRF (Mildenhall et al., 2020) models a volumetric 3D scene as a continuous function by mapping a 3D coordinate $\mathbf{x}=(x,y,z)$ and a 2D viewing direction $\mathbf{d}=(\theta,\phi)$ to a color (RGB) $\mathbf{c}$ and a density ($\sigma$). This function $F_{\theta}:(\mathbf{x},\mathbf{d})\mapsto(\mathbf{c},\mathbf{\sigma})$ is often parameterized by a neural network, a voxel grid structure (Fridovich-Keil et al., 2022), or a hybrid representation to accelerate performance (Müller et al., 2022; Sun et al., 2022a, b). Given a NeRF model trained on a set of 2D images taken from various viewpoints of a target scene, the accumulated color $C(\mathbf{r})$ along an arbitrary camera ray $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ is calculated with the quadrature rule by volume rendering (Max, 1995):
(1) $C(\mathbf{r})=\sum_{k=1}^{K}T_{k}\bigl{(}1-\exp({-\sigma_{k}\delta_{k}})\bigr{)}\mathbf{c}_{k},T_{k}=\exp{\Bigl{(}-\sum_{j=1}^{k-1}\sigma_{j}\delta_{j}\Bigr{)}}$
@@
3.1.2. Conditional Diffusion Models.
Recent T2I diffusion models (Rombach et al., 2022; Podell et al., 2023) are built with a U-net architecture (Ronneberger et al., 2015) integrated with convolutional layers and attention blocks (Vaswani et al., 2017). Within the model, attention blocks play a crucial role in correlating text with relevant parts of the deep features during image generation. Our work uses an open-source latent diffusion model (Podell et al., 2023), which includes a CLIP text encoder (Radford et al., 2021) for text embedding. The cross-attention between contextual text embedding and the deep features of the denoising network is calculated as follows:
(2) $\texttt{Attn}(Q,K,V)=\texttt{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}V\right)$
@@
(3) $\texttt{MultiHeadAttn}(Q,K,V)=\underset{j\in N_{h}}{\texttt{concat}}\Bigl{[}\texttt{Attn}(Q^{(j)},K^{(j)},V^{(j)})\Bigr{]}$
3.2. Style-NeRF2NeRF
@@
3.2.1. Style-Aligned Image-to-Image Generation.
Given a set of source view images $\{I^{(i)}\}(i=1,\ldots,N)$, our first goal is to generate a corresponding set of stylized view images $I^{(i)}_{c}=U_{\theta}(I^{(i)},c)$ under a text condition $c$ with as much perceptual view consistencies among images where $U_{\theta}$ consists of a sampling process such as DDIM (Song et al., 2020a).
Although T2I diffusion models can generate rich images with arbitrary text prompts, merely sharing the same prompt across different source views is insufficient to generate stylized images with a perceptually consistent style. To alleviate this problem, we apply a fully-shared-attention variant of a style-aligned image generation method proposed by (Hertz et al., 2023b). Let $Q_{i},K_{i},V_{i}$ be the queries, keys, and values from a deep feature $\phi_{i}$ for view $I^{(i)}$, then we generate $n$ stylized views simultaneously using the following fully-shared-attention:
(4) $\texttt{Attn}(Q_{i},K_{1\ldots n},V_{1\ldots n})\$
(5) $K_{1\ldots n}=[K_{1},K_{2},\ldots K_{n}]^{T},V_{1\ldots n}=[V_{1},V_{2},\ldots V_{n}]^{T}$
@@
3.2.2. Conditioning on Source Views.
To further strengthen perceptual consistencies across multi-view frames, we attach a depth-conditioned ControlNet (Zhang et al., 2023) and optionally enable SDEdit (Meng et al., 2021) for conditioning on the source view. As for the depth inputs, we may either render the corresponding depth maps from the source NeRF or use an off-the-shelf depth estimator model such as MiDaS (Ranftl et al., 2020).
@@
3.2.3. NeRF Fine-Tuning.
@@
Although the stylized multi-view images are a good starting point for fine-tuning the source NeRF, we found that using a common RGB pixel loss is prone to over-fitting due to ambiguities in 3D geometry and color. Therefore, an alternative loss function that reflects the perceptual similarity is preferred for guiding the 3D style-transfer process. To meet our requirement, we employ the Sliced Wasserstein Distance loss (SWD loss) (Heitz et al., 2021).
3.3. Sliced Wasserstein Distance Loss.
Feature statistics of pre-trained Convolutional Neural Networks (CNNs) such as VGG-19 (Simonyan and Zisserman, 2014) are known to be useful for representing a style of an image (Gatys et al., 2015; Johnson et al., 2016; Huang and Belongie, 2017; Li et al., 2017; Luan et al., 2017). In our study we employ the SWD loss originally proposed for texture synthesis (Heitz et al., 2021) as the loss term to guide the style-transfer process for NeRF.
@@
(7) $\mathcal{L}_{style}=\sum_{l=1}^{L}\mathcal{L}_{SWD}(p^{l},\hat{p}^{l})$
@@
(8) $\mathcal{L}_{SWD}=\sum_{l=1}^{L}\mathbb{E}_{V}[\mathcal{L}_{SW1D}(p^{l}_{V},\hat{p}^{l}_{V})]$
(9) $\mathcal{L}_{SW1D}(p^{l}_{V},\hat{p}^{l}_{V})=\frac{1}{|p^{l}_{V}|}\lVert\texttt{sort}(p^{l}_{V})-\texttt{sort}(\hat{p}^{l}_{V})\rVert^{2}$
Expectation over random projections $V$ provides a good approximation in practice and an optimized distribution is proven to converge to the target distribution. SWD is known to capture the complete target distribution (Pitie et al., 2005) as described below:
(10) $\mathcal{L}_{SW}(I,\hat{I})=0\implies p^{l}=\hat{p}^{l},\forall l\in{1,\ldots,L}$
@@
Figure 3. Sliced Wasserstein Distance. $p$ and $\hat{p}$ are projected onto a random unit direction $V$ (left). The $1$-dimensional Wasserstein distance can be calculated by taking the $L^{2}$ difference between the sorted projections $p$ and $\hat{p}$ (right). Expectation over random $V$ vectors is a practical approximation of the $N$-dimensional Wasserstein distance.
3.4. Style Blending.
@@
(11) $\mathcal{L}_{style}(I_{1},I_{2},\hat{I})=\sum_{l=1}^{L}\left(t\mathcal{L}_{SWD}(p_{1}^{l},\hat{p}^{l})+(1-t)\mathcal{L}_{SWD}(p_{2}^{l},\hat{p}^{l})\right)$
@@
Figure 4. Style Interpolation. An example of style blending using the Wasserstein barycenter between two different style prompts ”A person like Marilyn Monroe, pop art style” and ”A person like Steve Jobs”.
3.5. Implementation Details.
@@
We employ Stable Diffusion XL (Podell et al., 2023) as a backbone for the style-aligned image-to-image diffusion pipeline. For NeRF representation, we use the ”nerfacto” model implemented in NeRFStudio (Tancik et al., 2023). Due to memory constraints, we generate up to $18$ views simultaneously with a fixed seed across all $N$ source views. In our experiments, we generated the target images with $50$ denoising steps using a range of classifier-free guidance weights (mostly between 5 and 30) depending on the scene or the style text. While we use depth maps rendered by NeRF for relatively compact and forward-facing scenes, we opt for depth estimations from the MiDaS model (Ranftl et al., 2020) for large-scale outdoor scenes.
@@
As the image editing is performed before NeRF training, our method allows users to test with different text prompts and parameters (e.g., text guidance scale, SDEdit strength) beforehand. Additionally, our straight-forward NeRF training without Iterative DU (Haque et al., 2023) or score distillation sampling (SDS) loss (Poole et al., 2022) allows the training process to run with less GPU memory as editing by a diffusion model is not necessary during NeRF updates. We also verify the importance of style-alignment in the ablation study. Please refer to the supplementary material for more implementation details. The overall pipeline of our method is shown in Figure 2.
@@
Figure 5. Effect of Style-Alignment. An example of source view conversion applied to ”Bear” scene using a text prompt ”A water painting of a brown bear” with and without shared-attention mechanism within the diffusion pipeline. We find that a fully-shared-attention variant of the style-aligned diffusion model (Hertz et al., 2023b) greatly improves style consistencies among generated views.
4. Results
We run our experiments on several real-world scenes, including the Instruct-NeRF2NeRF (Haque et al., 2023) dataset captured by a smartphone and a mirrorless camera with camera poses extracted by COLMAP (Schönberger and Frahm, 2016) and PolyCam (pol, [n. d.]). The dataset contains large-scale 360 scenes, objects, and forward-facing human portraits.